# Stochastic Spiking Neural Networks with First-to-Spike Coding

### v1

@@ I Introduction
Although the vast majority of algorithm development and applications in SNNs employ deterministic neuron models such as the deterministic Spike Response Model (SRM) [16], Integrate and Fire (IF), and Leaky Integrate and Fire (LIF) models [17], it is important to recognize that biological neurons generate spikes in a stochastic fashion [18]. Furthermore, deterministic neuron models are discontinuous and non-differentiable, which presents substantial challenges in the application of gradient-based optimization methods. On the other hand, stochastic neuron models smoothen the network model to be continuously differentiable [19] and therefore have the potential to offer enhanced efficiency and robustness.
Recently, there has been growing interest in exploiting stochastic devices in the neuromorphic hardware community [20], [21]. With the scaling of device dimensions, memristive devices lose their programming resolution and are characterized by increased cycle-to-cycle variation. Work has started in earnest to design stochastic state-compressed SNNs using such scaled neuromorphic devices that exhibit iso-accuracies in comparison to their multi-bit deterministic counterparts enabled by the alternate encoding of information in the probability domain [22, 23]. While there is significant progress in the domain of stochastic SNN training algorithms, there remains a noticeable gap in the design of SNN architectures that integrate the benefits of both stochastic neuronal computing and temporal information coding. Most of the current efforts on directly training stochastic SNNs with temporal coding have primarily demonstrated success on simple datasets and shallow network structures [24, 25, 26]. Scaling these networks to deeper architectures and more complex datasets presents significant challenges. Further, as we demonstrate in this work, many of the supposed benefits of temporal encoding, like enhanced spiking sparsity, may not necessarily hold true for deep architectures. This necessitates a co-design approach to identify the relative tradeoffs of stochastic temporally encoded SNNs in terms of accuracy, latency, sparsity, energy cost, and robustness.
@@ III-B Stochastic SNN
Contrary to the usage of SG for deterministic neuron models, another way to solve the discontinuous spiking nonlinearity is by using stochastic neuron models. Inspired by [24], we integrate stochastic LIF neurons with First-to-Spike coding. The membrane potential is computed by using the following equation:
@@ IV Results
In this section, we evaluate the accuracy, latency, sparsity, energy cost, and noise sensitivity of different types of models to evaluate the influence of information encoding, computing scheme and training methods independently: ANN-SNN conversion utilizing deterministic neurons and rate coding (D-R-CONV) [38], BPTT trained models with deterministic neurons utilizing rate coding (D-R-BPTT) [11], deterministic neural networks trained by BPTT utilizing First-to-Spike coding (D-F-BPTT) and the stochastic neural networks trained by BPTT utilizing First-to-Spike coding (S-F-BPTT). Acronyms are used to simplify the naming that reflects its key features: The first part indicates the type of neurons: deterministic (D) and stochastic (S), the second part denotes the coding method: rate coding (R) and First-to-Spike coding (F), and the third part represents the training method: ANN-SNN conversion (CONV) and Backpropagation Through Time (BPTT). We will use these acronyms throughout the remainder of the paper for brevity. We conduct experiments for three neural network architectures, ranging from shallow to deep: 2-layer MLP, LeNet5, and VGG15.
@@ IV-C Quantitative Analysis
Inference Latency: In SNNs, reducing latency without sacrificing accuracy can be a critical goal, allowing for faster and more energy-efficient computation. In particular, for the D-R-CONV and D-R-BPTT models, the optimal number of timesteps is determined by identifying the saturation point on a plot of timesteps versus accuracy, where further increase in timesteps no longer significantly improves model accuracy. The differences in SNN inference latency in terms of timesteps are noted in Table II. Compared to other temporal coding models, First-to-Spike coding models show a significantly lower latency. In the First-to-Spike coding scheme, the result is based on which neuron in the output layer is the first to generate a spike. This approach requires only a single spike in the output layer to ascertain the result, reducing the number of timesteps significantly. On the other hand, rate coding relies on the frequency of spikes over time, and therefore the network needs a longer observation window to establish an accurate spike rate. This phenomenon is magnified when the dataset becomes complex. On the CIFAR-10 dataset, the rate coding approaches require a substantially higher number of timesteps to achieve the same level of accuracy compared to the models employing First-to-Spike coding. Interestingly, we find that the S-F-BPTT model reduces the latency even further in comparison to the D-F-BPTT model. The stochastic nature of spike generation in S-F-BPTT models may cause an output spike generation even when the input stimulus is not too strong or the membrane potential is low, allowing for a faster response to the input.
Sparsity: Spiking sparsity in SNNs is an important metric for evaluating the efficiency and functionality of models. The average spiking rate of a particular layer, defined as the average number of spikes that a neuron generates over a fixed time interval, is utilized to quantitatively represent sparsity. In this context, a higher average spike rate indicates lower sparsity, and vice versa. The average spiking rate of each model across various layers for the LeNet5 architecture trained on the MNIST dataset and VGG15 architecture trained on the CIFAR-10 dataset is shown in Fig. 1. Contrary to the common assumption of higher sparsity in temporal coding models than in rate coding models, the figure shows that first-to-spike models do show higher sparsity in the final layer, with the hidden layers presenting a contrary trend. The main reason is the necessity to encode the same information with reduced latency for temporally encoded models, which demands an increased spike count. It also explains why the S-F-BPTT model has the highest spiking rate with the lowest latency. Moreover, to achieve a reliable and consistent output in the presence of stochasticity, the stochastic SNN model needs to increase its spiking rate. This compensates for the unpredictability of individual spikes, ensuring that the overall signal transmission between neurons remains stable and correct.
@@
Figure 1: Average spiking rate of different neuronal layers for different SNN models (S-F-BPTT, D-F-BPTT, D-R-BPTT, D-R-CONV) corresponding to (a) LeNet5 on the MNIST dataset (b) VGG15 on the CIFAR-10 dataset.
@@ Acknowledgments
This material is based upon work supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number #DE-SC0021562, the National Science Foundation grants CCSS #2333881, CCF #1955815, CAREER #2337646 and EFRI BRAID #2318101 and by Oracle Cloud credits and related resources provided by the Oracle for Research program.

### v3

@@ I Introduction
Although the vast majority of algorithm development and applications in SNNs employ deterministic neuron models such as the deterministic Spike Response Model (SRM) [16], Integrate and Fire (IF), and Leaky Integrate and Fire (LIF) models [17], it is important to recognize that biological neurons generate spikes in a stochastic fashion [18]. Furthermore, deterministic neuron models are discontinuous and non-differentiable, which presents substantial challenges in the application of gradient-based optimization methods. On the other hand, stochastic neuron models smoothen the network model to be continuously differentiable [19] and therefore have the potential to offer enhanced efficiency and robustness. Additionally, stochasticity can enhance generalization performance by improving fault tolerance [20] and by preventing overfitting [21].
Recently, there has been growing interest in exploiting stochastic devices in the neuromorphic hardware community [22], [23]. With the scaling of device dimensions, memristive devices lose their programming resolution and are characterized by increased cycle-to-cycle variation. Work has started in earnest to design stochastic state-compressed SNNs using such scaled neuromorphic devices that exhibit iso-accuracies in comparison to their multi-bit deterministic counterparts enabled by the alternate encoding of information in the probability domain [24, 25]. Ref. [26] proposes noisy spiking neural network (NSNN) which leverages stochastic noise as a resource for training SNNs using rate-based coding. While there is significant progress in the domain of stochastic SNN training algorithms, there remains a noticeable gap in the design of SNN architectures that integrate the benefits of both stochastic neuronal computing and temporal information coding. Most of the current efforts on directly training stochastic SNNs with temporal coding have primarily demonstrated success on simple datasets and shallow network structures [27, 28, 29]. Scaling these networks to deeper architectures and more complex datasets presents significant challenges. Further, as we demonstrate in this work, many of the supposed benefits of temporal encoding, like enhanced spiking sparsity, may not necessarily hold true for deep architectures. This necessitates a co-design approach to identify the relative tradeoffs of stochastic temporally encoded SNNs in terms of accuracy, latency, sparsity, energy cost, and robustness.
@@ III-B Stochastic SNN
On the other hand, the introduction of stochasticity can efficiently smoothen out discontinuous spiking nonlinearities. Inspired by [27], we integrate stochastic LIF neurons with First-to-Spike coding. The membrane potential is computed by using the following equation:
@@ IV Results
In this section, we evaluate the accuracy, latency, sparsity, energy cost, and noise sensitivity of different types of models to evaluate the influence of information encoding, computing scheme, and training methods independently: ANN-SNN conversion utilizing deterministic neurons and rate coding (D-R-CONV) [41], BPTT trained models with deterministic neurons utilizing rate coding (D-R-BPTT) [11], deterministic neural networks trained by BPTT utilizing First-to-Spike coding (D-F-BPTT) and the stochastic neural networks trained by BPTT utilizing First-to-Spike coding (S-F-BPTT). It is worth noting that First-to-Spike coding is only applied in the final output layer to leverage the spike timing at which a neuron first spikes to encode information. Acronyms are used to simplify the naming that reflects its key features: The first part indicates the type of neurons: deterministic (D) and stochastic (S), the second part denotes the coding method: rate coding (R) and First-to-Spike coding (F), and the third part represents the training method: ANN-SNN conversion (CONV) and Backpropagation Through Time (BPTT). We will use these acronyms throughout the remainder of the paper for brevity. We conduct experiments for three neural network architectures, ranging from shallow to deep: 2-layer MLP, LeNet5, and VGG15.
@@ IV-C Quantitative Analysis
Inference Latency: In SNNs, reducing latency without sacrificing accuracy can be a critical goal, allowing for faster and more energy-efficient computation. In particular, for the D-R-CONV and D-R-BPTT models, the optimal number of timesteps is determined by identifying the saturation point on a plot of timesteps versus accuracy, where further increase in timesteps no longer significantly improves model accuracy. For the S-F-BPTT and D-F-BPTT models, the inference latency is determined by averaging the number of timesteps at which the first spike is detected in the output layer across all input data. The differences in SNN inference latency in terms of timesteps are noted in Table II. Compared to other temporal coding models, First-to-Spike coding models show a significantly lower latency. The First-to-Spike coding scheme requires only a single spike in the output layer to ascertain the result, reducing the number of timesteps significantly. On the other hand, rate coding relies on the frequency of spikes over time, and therefore the network needs a longer observation window to establish an accurate spike rate. This phenomenon is magnified when the dataset becomes complex. On the CIFAR-10 dataset, the rate coding approaches require a substantially higher number of timesteps to achieve the same level of accuracy compared to the models employing First-to-Spike coding. Interestingly, we find that the S-F-BPTT model reduces the latency even further in comparison to the D-F-BPTT model. The stochastic nature of spike generation in S-F-BPTT models may cause an output spike generation even when the input stimulus is not too strong or the membrane potential is low, allowing for a faster response to the input.
@@
Figure 1: Average spiking rate of different neuronal layers for different SNN models (S-F-BPTT, D-F-BPTT, D-R-BPTT, D-R-CONV) corresponding to (a) LeNet5 on the MNIST dataset, (b) VGG15 on the CIFAR-10 dataset.
Sparsity: Spiking sparsity in SNNs is an important metric for evaluating the efficiency and functionality of models. The average spiking rate of a particular layer, defined as the average number of spikes that a neuron generates over a fixed time interval, is utilized to quantitatively represent sparsity. In this context, a higher average spike rate indicates lower sparsity, and vice versa. The average spiking rate of each model across various layers for the LeNet5 architecture trained on the MNIST dataset and VGG15 architecture trained on the CIFAR-10 dataset is shown in Fig. 1. Contrary to the common assumption of higher sparsity in temporal coding models than in rate coding models, the figure shows that first-to-spike models do show higher sparsity in the final layer, with the hidden layers presenting a contrary trend, especially for the stochastic model, as shown in the results. The main reason is the necessity to encode the same information with reduced latency for temporally encoded models, which demands an increased spike count. It also explains why the S-F-BPTT model has the highest spiking rate with the lowest latency. Moreover, to achieve a reliable and consistent output in the presence of stochasticity, the stochastic SNN model needs to increase its spiking rate. This compensates for the unpredictability of individual spikes, ensuring that the overall signal transmission between neurons remains stable and correct.
@@ Acknowledgments
This material is based upon work supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number #DE-SC0021562, the U.S. National Science Foundation under award No. CCSS #2333881, CCF #1955815, CAREER #2337646 and EFRI BRAID #2318101 and by Oracle Cloud credits and related resources provided by the Oracle for Research program.